# Data-adaptive structural change-point detection via isolation

### v1

@@ 3 Theory
DAIS requires knowledge of the standard deviation, $\sigma$, of the data generation mechanism in (1). If $\sigma$ is unknown, it can be estimated using the Median Absolute Deviation method (MAD), with the estimator defined as $\hat{\sigma}:=C\times\text{median}(\left|\boldsymbol{x}-\text{median}(\boldsymbol{x})\right|)$ for $\boldsymbol{x}=(x_{1},x_{2},\dots x_{T})$, as proposed by [21]. In the case of Gaussian data, for $C=1.4826$ it is a consistent estimator of the population standard deviation $\sigma$ as proven by [36]. In this section, we assume that $\epsilon_{t}$ in (1) are i.i.d. random variables following the $\mathcal{N}(0,1)$ distribution. This assumption is commonly used in the literature to prove the theoretical results. We also assume that $\sigma=1$ in order to make notation easier in the proof of the theorems presented in this section. With this assumption, the model becomes
@@
In Sections 3.1 and 3.2 we provide the theoretical results on the consistency of the algorithm for the cases of piecewise-constant and continuous, piecewise-linear signals, respectively. The proofs of Theorems 1 and 2 can be found in Appendix B and D. In Section 3.3, we return to the discussion of Section 2.3 and present some results in this specific case where $\epsilon_{t}$ in (1) are i.i.d. from the $\mathcal{N}(0,1)$ distribution.
@@ 3.1 Piecewise-constant signals
$C_{s,e}^{b}\left(\textbf{X}\right)=\left|\tilde{X}^{b}_{s,e}\right|=\left|\sqrt{\frac{e-b}{\ell(b-s+1)}}\sum^{b}_{t=s}X_{t}-\sqrt{\frac{b-s+1}{\ell(e-b)}}\sum^{e}_{t=b+1}X_{t}\right|,$ (9)
where $s\leq b<e$ and $\ell=e-s+1$. The value of $C_{s,e}^{b}\left(\textbf{X}\right)$ is small if $b$ is not a change-point and large otherwise.
@@
    The location of the largest difference $d_{s,e}$ as defined in (2), satisfies $d_{s,e}\in\bigcup_{j=1}^{N}[r_{j}-\frac{\delta_{j-1}}{2}+\frac{3\delta_{T}}{4n},r_{j}+\frac{\delta_{j}}{2}-\frac{3\delta_{T}}{4n}]$ every time the algorithm is applied to an interval $[s,e]$ which contains at least one change-point.
@@
    The minimum distance, $\delta_{T}$, between two successive change-points and the minimum magnitude of jumps $\underline{f}_{T}$ are connected by $\sqrt{\delta_{T}}\underline{f}_{T}\geq\underline{C}\sqrt{\log T}$ for a large enough constant $\underline{C}$.
The first assumption is a formal statement of what is discussed in Section 2.3. It is not a strict assumption and the exact probability of the consecutive pairwise difference being larger at a change-point compared to a point where no change occurs, in the case where $\epsilon_{t}$ is assumed to be Gaussian, is provided in Section 3.3. Below, we provide the relevant theorem for the consistency of DAIS in accurately estimating the true number and the locations of the change-points in the case of piecewise-constant signals.
@@
The lower bound of the probability in (10) is $1-\mathcal{O}(1/T)$. From Theorem 1, we can conclude that in order to be able to match the estimated change-points with the true ones, $\delta_{T}$ needs to be larger than $\max_{j=1,2,\ldots,N}\left|\hat{r}_{j}-r_{j}\right|$, which means that $\delta_{T}$ must be at least of order $\mathcal{O}(\log T)$. Also, assumption (A2) ensures that the rate attained for $\delta_{T}\underline{f}_{T}^{2}$, which characterizes the complexity of the problem, is $\mathcal{O}(\log T)$ and, as [9] argues, the lowest possible $\delta_{T}\underline{f}_{T}^{2}$ that allows detection of the change-points is $\mathcal{O}(\log T-\log\log T)$. Therefore, the rate obtained by DAIS is optimal up to a double logarithmic factor.
@@ 3.2 Continuous, piecewise-linear signals
For the case of continuous, piecewise-linear signals $f_{t}=\mu_{j,1}+\mu_{j,2}t$ for $t=r_{j-1}+1,\ldots,r_{j}$ and $f_{r_{j}-1}+f_{r_{j}+1}\neq 2f_{r_{j}}$ with the additional constraint $\mu_{j,1}+\mu_{j,2}r_{j}=\mu_{j+1,1}+\mu_{j+1,2}r_{j}$ so that the signal is continuous. [5] shows that the best choice of contrast function is $C_{s,e}^{b}\left(\textbf{X}\right)=\left|\langle\boldsymbol{X},\boldsymbol{\phi^{b}_{s,e}}\rangle\right|$, where the contrast vector $\boldsymbol{\phi^{b}_{s,e}}=(\phi^{b}_{s,e}(1),\dots,\phi^{b}_{s,e}(T))$ is given by
$\phi^{b}_{s,e}(t)=\left\{\begin{array}[]{lll}\alpha^{b}_{s,e}\beta^{b}_{s,e}\big{[}(e+2b-3s+2)t-(be+bs-2s^{2}+2s)\big{]},&t\in\{s,\dots,b\},\\ -\frac{\alpha^{b}_{s,e}}{\beta^{b}_{s,e}}\big{[}(3e-2b-s+2)t-(2e^{2}+2e-be-bs)\big{]},&t\in\{b+1,\ldots,e\},\\ 0,&\text{otherwise}\end{array}\right.$ (11)
@@
Under the Gaussianity assumption, $C_{s,e}^{b}\left(\textbf{X}\right)$ as defined above, is maximized at the same point as the generalized log-likelihood ratio for all possible single change-points within $[s,e)$.
Before presenting the theoretical result for the consistency of the number and location of the estimated change-points in the case of continuous piecewise-linear signals, we require the following assumption, which is equivalent to (A2) in Section 3.1, but now under the current setting of changes in the slope.
@@
Again, The lower bound of the probability is $1-\mathcal{O}(1/T)$ and the difficulty of the problem, $\delta_{T}^{3/2}\underline{f}_{T}$ is analogous to $\sqrt{\delta_{T}}\underline{f}_{T}$ in the case of piecewise-constant signals. The proof of Theorem 2 can be found in Appendix D.
3.3 Probability of isolation
We now return to the discussion of Section 2.3. Under the assumption of $\epsilon_{t}$ being i.i.d. random variables following the $\mathcal{N}(0,1)$ distribution, and in the case that $f_{t}$ is piecewise-constant, we note that in order to calculate the probability that the largest difference is exactly at the location of a change-point, we need to calculate the following:
@@ 3.3 Probability of isolation
For $t\notin\{r_{j}-1,r_{j},r_{j}+1\},j\in\{1,\ldots,N\}$, we focus on the probability that the largest difference occurs at the location of one of the $N$ change-points. Note that $|X_{r_{j}+1}-X_{r_{j}}|$ and $|X_{t+1}-X_{t}|$ are independent for these values of $t$. We will work on
@@
which is the probability that the observations at the change-point have a larger absolute difference than a pair of observations in the signal where no change occurs. Using (1), the fact that $\epsilon_{t}\sim\mathcal{N}(0,1)$ and that $f_{t^{\prime}+1}-f_{t^{\prime}}=0$ for $t^{\prime}\neq r_{j},j\in\{1,\ldots,N\}$, (14) can be written as
@@
where $\alpha_{j}=(f_{r_{j}+1}-f_{r_{j}})/\sigma$ and $Z_{2},\tilde{Z}_{2}\sim\mathcal{N}(0,2)$ are independent. In the above equation, $|\alpha_{j}+Z_{2}|$ follows the Folded Normal (FN) distribution with mean $\alpha_{j}$ and variance 2. Similarly, $|\tilde{Z}_{2}|$, follows the same distribution with mean 0 and variance 2, which reduces to the simpler case of the Half Normal (HN) distribution with variance 2. So, (15) can be equivalently written as
$\mathbb{P}(Y_{1}\geq Y_{2})$ (16)
for $Y_{1}\sim FN(\alpha_{j},2)$ and $Y_{2}\sim HN(0,2)$. Denoting by $f_{Y_{1}}$ the probability density function of $Y_{1}$ and $F_{Y_{2}}$ the cumulative density function of $Y_{2}$, (16) can be written as
@@
where $t\textrm{erf}(z)=\frac{2}{\sqrt{\pi}}\int_{0}^{z}e^{-t^{2}}\rm{d}$t is the Gauss error function. Using [32] (equation (38) on page 10), we can calculate the above in terms of $\alpha_{j}$ as
@@
$\mathbb{P}(Y_{1}\geq Y_{2})=\frac{1}{2}\biggl{[}1+\textrm{erf}\Bigl{(}\frac{\alpha_{j}}{2\sqrt{2}}\Bigl{)}^{2}\biggr{]}.$
@@
It is easy to see now that this probability goes to 1 as $\alpha_{j}$ gets larger and in fact $\mathbb{P}(Y_{1}\geq Y_{2})>0.99$ for $\alpha_{j}=5.15$. However, calculating or bounding from below the probability in (13) is a very difficult task due to the fact that $|X_{t}-X_{t-1}|$ and $|X_{t-1}-X_{t-2}|$ are dependent. Therefore, we provide results based on simulations.
We ran a simulation study to observe the behavior of the probability of the largest difference being in a neighborhood that guarantees detection in an interval where the change-point is isolated as $\alpha=\min_{j=1,2,\ldots,N}{\alpha_{j}}$ increases. By the definition of $\alpha$, it holds that $\alpha=\underline{f}_{T}/\sigma$, where $\underline{f}_{T}=\min_{j=0,1,\ldots,N}\left|f_{r_{j+1}}-f_{r_{j}}\right|$. The location of the largest difference is close to the change-point with the largest jump in the interval under consideration, and so in the case of multiple change-points, the probability in question depends on the value of the largest magnitude of change. In these simulations we use the value of $\underline{f}_{T}$, to obtain the smallest probability of the largest difference being in a location that guarantees detection throughout the whole algorithm, until all change-points have been detected.
@@
For the simulation study, we used one change-point, $r$, with $\delta_{T}$ fixed at value 20 and set $\sigma=1$. We repeated this 10000 times for various values of the length $T$. We calculated the Monte Carlo estimates of the probabilities that the location of the largest difference ensured that detection occurred in an interval where the change-point is isolated as $T$ increased. In other words, using (7), we calculated the proportion of times that
$d_{1,T}\in\Bigl{[}\frac{r+1}{2}+\frac{3\delta_{T}}{4n},\frac{T+r}{2}-\frac{3\delta_{T}}{4n}\Bigr{]},$
@@
where $n$ is chosen as $n=\frac{\delta_{T}}{2\lambda_{T}}$ for $\lambda_{T}=3$. In order to be consistent with assumption (A2), $\underline{f}_{T}=c\sqrt{\log T}$, where $c>0$ is a constant. In our simulation study, $c$ is chosen such that $\alpha$ is equal to five when $T=50$. This choice of $c$ led to an early convergence of the relevant proportion; please, see Table 1. Smaller values of $c$ would be acceptable but larger values of the sample size, $T$, would be required until convergence to one becomes apparent. The results in Table 1 show that the probability of the largest difference occurring at a point that guarantees isolation in the interval where detection occurred increases with the length of the signal. This is expected as $\underline{f}_{T}=\mathcal{O}(\sqrt{\log(T)})$.
@@
Table 1: Proportion of times that the location of the largest difference was in an interval where detection while the change-point is isolated is guaranteed.

### v2

@@ 3 Theory
DAIS requires knowledge of the standard deviation, $\sigma$, of the data generation mechanism in (1). If $\sigma$ is unknown, it can be estimated using the Median Absolute Deviation method (MAD), with the estimator defined as $\hat{\sigma}:=C\times\text{median}(\left|\boldsymbol{x}-\text{median}(\boldsymbol{x})\right|)$ for $\boldsymbol{x}=(x_{1},x_{2},\dots x_{T})$, as proposed by [26]. In the case of Gaussian data, for $C=1.4826$, $\hat{\sigma}$ is a consistent estimator of the population standard deviation $\sigma$ as proven by [43]. In this section, we assume that $\epsilon_{t}$ in (1) are i.i.d. random variables following the $\mathcal{N}(0,1)$ distribution. This assumption is commonly used in the literature to prove theoretical consistency results. Without loss of generality, we let $\sigma=1$ in order to make notation easier in the proof of the theorems presented in this section. The model becomes
@@
In Sections 3.1 and 3.2 we provide the theoretical results on the consistency of the algorithm for the cases of piecewise-constant and continuous, piecewise-linear signals, respectively. The proofs of Theorems 1 and 2 can be found in Appendix C and D, respectively. In Section 6.3 we also provide theoretical results for the consistency in the case of multivariate signals and the outline of the proof can be found in Appendix E. In Section 3.3, we return to the discussion of Section 2.3 and present some results in the case when (7) holds, and $\epsilon_{t}$ in (1) are i.i.d. from the $\mathcal{N}(0,1)$ distribution.
@@ 3.1 Piecewise-constant signals
$C_{s,e}^{b}\left(\boldsymbol{X}\right)=\left|\sqrt{\frac{e-b}{\ell(b-s+1)}}\sum^{b}_{t=s}X_{t}-\sqrt{\frac{b-s+1}{\ell(e-b)}}\sum^{e}_{t=b+1}X_{t}\right|,$ (10)
where $s\leq b<e$ and $\ell=e-s+1$. The value of $C_{s,e}^{b}\left(\boldsymbol{X}\right)$ is small if $b$ is not a change-point and large otherwise.
@@
    The location of the largest difference $d_{s,e}$ as defined in (2), satisfies (7) every time the algorithm is applied to an interval $[s,e]$ which contains at least one change-point.
@@
    The minimum distance, $\delta_{T}$, between two successive change-points and the minimum magnitude of jumps $\underline{f}_{T}=\min_{j}\left|f_{r_{j}+1}-f_{r_{j}}\right|$ are connected by $\sqrt{\delta_{T}}\underline{f}_{T}\geq\underline{C}\sqrt{\log T}$ for a large enough constant $\underline{C}$.
The first assumption is a formal statement of what is discussed in Section 2.3 and is required to ensure that every change-point can be detected in an interval in which it is isolated. It is not a strict assumption as for appropriately chosen values of $n$ and $\lambda_{T}$, (7) includes any point besides the midpoint of two consecutive change-points. The exact probability of the consecutive pairwise difference being larger in absolute value at a change-point compared to a point where no change occurs, in the case where $\epsilon_{t}$ is assumed to be Gaussian, is provided in Section 3.3. (A2) is not a very restrictive assumption, for example if $\delta_{T}\geq O(\log T)$, then $\underline{f}_{T}$ is allowed to decrease as $T$ increases. Further discussion for (A2) is given after Theorem 1.
Below, we provide the relevant theorem for the consistency of DAIS in accurately estimating the true number and the locations of the change-points in the case of piecewise-constant signals.
@@
The lower bound of the probability in (11) is $1-\mathcal{O}(1/T)$. From Theorem 1, we can conclude that in order to be able to match the estimated change-points with the true ones, $\delta_{T}$ needs to be larger than $\max_{j=1,2,\ldots,N}\left|\hat{r}_{j}-r_{j}\right|$, which means that $\delta_{T}$ must be at least of order $\mathcal{O}(\log T)$. Assumption (A2) ensures that the rate attained for $\delta_{T}\underline{f}_{T}^{2}$, which characterizes the complexity of the problem, is $\mathcal{O}(\log T)$ and, as [11] argues, the lowest possible $\delta_{T}\underline{f}_{T}^{2}$ that allows detection of the change-points, in the case of piecewise-constant signals is $\mathcal{O}(\log T-\log\log T)$. Therefore, the rate obtained by DAIS is optimal up to a rather negligible double logarithmic factor. As mentioned in [53], the rate we obtain has been established in various different papers. Rates for other methods which are the same or comparable to ours can be found in [17].
@@ 3.2 Continuous, piecewise-linear signals
For the case of continuous, piecewise-linear signals, $f_{t}=\mu_{j,1}+\mu_{j,2}t$ for $t=r_{j-1}+1,\ldots,r_{j}$ and $f_{r_{j}-1}+f_{r_{j}+1}\neq 2f_{r_{j}}$ with the additional constraint $\mu_{j,1}+\mu_{j,2}r_{j}=\mu_{j+1,1}+\mu_{j+1,2}r_{j}$ so that the signal is continuous. Through a log-likelihood approach, [6] shows that contrast function is $C_{s,e}^{b}\left(\boldsymbol{X}\right)=\left|\langle\boldsymbol{X},\boldsymbol{\phi^{b}_{s,e}}\rangle\right|$, where the contrast vector $\boldsymbol{\phi^{b}_{s,e}}=(\phi^{b}_{s,e}(1),\dots,\phi^{b}_{s,e}(T))$ is given by
$\phi^{b}_{s,e}(t)=\left\{\begin{array}[]{lll}\alpha^{b}_{s,e}\beta^{b}_{s,e}\big{[}(e+2b-3s+2)t-(be+bs-2s^{2}+2s)\big{]},&t\in\{s,\dots,b\},\\ -\frac{\alpha^{b}_{s,e}}{\beta^{b}_{s,e}}\big{[}(3e-2b-s+2)t-(2e^{2}+2e-be-bs)\big{]},&t\in\{b+1,\ldots,e\},\\ 0,&\text{otherwise,}\end{array}\right.$ (12)
@@
Under the Gaussianity assumption, $C_{s,e}^{b}\left(\boldsymbol{X}\right)$, as defined above, is maximized at the same point as the generalized log-likelihood ratio for all possible single change-points within $[s,e)$.
Before presenting the theoretical result for the consistency of the number and location of the estimated change-points by DAIS in the case of continuous piecewise-linear signals, we require the following assumption, which is equivalent to (A2) in Section 3.1, but now under the current setting of changes in the slope.
@@
The proof of Theorem 2 can be found in Appendix D. Again, the lower bound of the probability is $1-\mathcal{O}(1/T)$ and the difficulty of the problem, $\delta_{T}^{3/2}\underline{f}_{T}$ is analogous to $\sqrt{\delta_{T}}\underline{f}_{T}$, which appears in (A2) for the case of piecewise-constant signals. For $\underline{f}_{T}\sim T^{1/2}$, the change-point detection accuracy is $O\left(T^{-1/3}\left(\log T\right)^{1/3}\right)$, which differs from the optimal rate $O(T^{-1/3})$ derived by [41] only by the logarithmic factor. This means that DAIS’s change-point detection rate is almost optimal.
3.3 Probability of guaranteed isolation of the change-point at detection
We now return to the discussion of Section 2.3, that concerns the connection between the location of the largest difference and the guarantee that the detection of the change-point closest to it occurs in an interval in which it is isolated. The probability of this happening is given by
$\mathbb{P}\left(d_{s,e}\in\left[1,r_{1}+\frac{\tilde{\delta}_{1}}{2}-\frac{3\delta_{T}}{4n}\right]\cup\left[r_{N}-\frac{\tilde{\delta}_{N-1}}{2}+\frac{3\delta_{T}}{4n},T\right]\cup\left\{\bigcup_{j=2}^{N-1}B_{n;j}\right\}\right)$ (14)
where $B_{n;j}$ is defined in (8). Under the assumption of $\epsilon_{t}$ being i.i.d. random variables following the $\mathcal{N}(0,1)$ distribution, and in the case that $f_{t}$ is piecewise-constant, we note that in order to calculate the probability that the largest difference is exactly at the location of a change-point, we need to calculate the following:
@@ 3.3 Probability of guaranteed isolation of the change-point at detection
[…]
@@
which is the probability that the observations at the change-point $r_{j}$, for any $j\in\{1,\ldots,N\}$, have a larger absolute difference than a pair of observations in the signal where no change occurs. Using (1), the fact that $\epsilon_{t}\sim\mathcal{N}(0,1)$, and that $f_{t^{\prime}+1}-f_{t^{\prime}}=0$ for $t^{\prime}\neq r_{j},j\in\{1,\ldots,N\}$, (16) can be written as
@@
where $\alpha_{j}=(f_{r_{j}+1}-f_{r_{j}})/\sigma$ and $Z_{2},\tilde{Z}_{2}\sim\mathcal{N}(0,2)$ are independent. In the above equation, $|\alpha_{j}+Z_{2}|$ follows the Folded Normal (FN) distribution with mean $\alpha_{j}$ and variance 2. Similarly, $|\tilde{Z}_{2}|$, follows the same distribution with mean 0 and variance 2, which reduces to the simpler case of the Half Normal (HN) distribution with variance 2. So, (17) can be equivalently written as $\mathbb{P}(Y_{1}\geq Y_{2})$ for $Y_{1}\sim FN(\alpha_{j},2)$ and $Y_{2}\sim HN(0,2)$. Denoting by $f_{Y_{1}}$ the probability density function of $Y_{1}$ and $F_{Y_{2}}$ the cumulative density function of $Y_{2}$, (17) can be written as
@@
$=\ \mathbb{E}_{Y_{1}}\Bigl{[}F_{Y_{2}}(Y_{1})\Bigr{]}=\ \mathbb{E}_{Y_{1}}\Bigl{[}\textrm{erf}\Bigl{(}\frac{Y_{1}}{2}\Bigr{)}\Bigr{]}$
@@
where $t\textrm{erf}(z)=\frac{2}{\sqrt{\pi}}\int_{0}^{z}e^{-t^{2}}\rm{d}$t is the Gauss error function. Using [38] (equation (38) on p.10), we can calculate the above in terms of $\alpha_{j}$ as
@@
$\mathbb{P}(Y_{1}\geq Y_{2})=\frac{1}{2}\biggl{[}1+\textrm{erf}\Bigl{(}\frac{\alpha_{j}}{2\sqrt{2}}\Bigl{)}^{2}\biggr{]},$ (18)
@@
so we have an explicit expression for (17). It is easy to see now that this probability goes to 1 as $\alpha_{j}$ gets larger. However, calculating or bounding from below the probability in (15), and subsequently in (14), are very difficult tasks due to the fact that $|X_{t}-X_{t-1}|$ and $|X_{t-1}-X_{t-2}|$ are dependent. Therefore, we provide results based on simulations.
We want to note that when $N=1$, the interval in assumption (A1) becomes $[1,T]$, which implies that isolation of the change-point in the interval in which detection occurs is guaranteed for all possible values of the largest difference. Also, in a data sequence with $N\geq 2$ change-points, we expect the order that they are detected to be with decreasing magnitude of change. This means that the change-point with the smallest magnitude of change will be detected last, in an interval in which it is certainly isolated, so the second smallest magnitude of change in the data sequence is the one that needs to be ‘large enough’ to ensure that all change-points will be isolated in an interval where detection can occur. We performed some simulations to check what values are considered ‘large enough’.
@@
In our simulations we consider a data sequence with $\sigma=1$ and two change-points at locations $r_{1}=20$ and $r_{2}=T-20$, with equal magnitudes of change, $\alpha=\alpha_{1}=\alpha_{2}$. It is easy to see that $\delta_{T}=20$ and $\tilde{\delta}_{1}=T-40$. We calculated the Monte Carlo estimates of (14), which is the probability that the location of the largest difference ensured that detection occurred in an interval where the change-point is isolated as $T$ increased, using 10.000 repetitions for each value of $T$. In other words, choosing $n=\frac{\delta_{T}}{2\lambda_{T}}$ for $\lambda_{T}=3$ we calculated the proportion of times that $d_{1,T}\in\left[1,\frac{T}{2}-4.5\right]\cup\left[\frac{T}{2}+4.5,T\right].$
@@
In order to be consistent with assumption (A2), we have $\underline{f}_{T}=c\sqrt{\log T}$, where $c>0$ is a constant; in this case, $c=\frac{1}{2}$ was chosen for the sake of running the simulations and there is no other reason behind this specific choice. The results in Table 1 show that the probability of the largest difference occurring at a point that guarantees isolation in the interval where detection occurred increases with the length of the signal. This is expected as can be seen by (18) and using the fact that $\underline{f}_{T}=\mathcal{O}(\sqrt{\log(T)})$ if $\delta_{T}$ is constant, which is the case in this simulation study. It is worth noting that the proportions are high even for small $T$ and $\alpha$, which provides evidence that the probability of isolation is extremely high.
@@
Table 1: Proportion of times that the location of the largest difference was in an interval where detection while the change-points are isolated is guaranteed.